# Model for charge-carriers spectra in topological semimetals of TaAs family

### v1

@@ Abstract
We propose a four-bands model describing the electron energy spectra near the Weyl points in the topological semimetals of the TaAs family (TaAs, TaP, NbAs, NbP). This model takes into account the fact that these Weyl points result from the band-contact lines which would exist in the mirror-reflection planes of these materials if the spin-orbit interaction were absent in them. Within this model, we obtain conditions for the existence of the Weyl points, determine their positions in the Brillouin zone, and derive the explicit formula for dispersion of the bands along the straight line connecting the two close Weyl points with opposite topological charges. Using NbP as an example, the values of the parameters defining the model spectrum are found. The obtained results show that for the semimetals of the TaAs family, the charge-carriers spectrum in the vicinity of the two close Weyl points can be analyzed without complex band-structure calculations.
@@ I Introduction
where $p_{i}$ ($i=x,y,z$) are the components of the quasimomentum, the velocities $v_{i}$ specify the splitting of the bands at the Weyl point ${\bf p}=0$ that has the energy $\varepsilon^{W}$, and the constant vector ${\bf a}=(a_{x},a_{y},a_{z})$ with the dimension of the velocity determines the so-called tilt of the spectrum. Moreover, the use of the additional simplification ${\bf a}=0$ is not rare in publications. However, only pairs of Weyl points with the opposite topological charges (the Chern numbers) can arise in the Brillouin zone of crystals, and the distance between these points of each pair is relatively small. This means that the region of the Brillouin zone where the linear dispersion of the bands really occurs is small, and a deviation from the linear dispersion can manifest itself even for a small difference between the Fermi energy $\varepsilon_{F}$ and the energy $\varepsilon^{W}$ of the Weyl points in the pair. Therefore, in order to correctly find the Fermi surface and its extremal cross sections, calculate the Landau levels and various physical quantities for the Weyl semimetals in a magnetic field, it may be necessary to consider the energy spectrum of the pair of the Weyl points [3] (see also recent papers [4, 5, 6]). Moreover, if $|\varepsilon_{F}-\varepsilon^{W}|$ is larger than the energy barrier separating the points, for the correct description of the charge-carriers spectrum, it is necessary to take into account four close energy bands that exist near the Weyl points. These four bands result from the nodal line (lying in the mirror-reflection plane) which would exist in these semimetals if the spin-orbit were absent in them [7]. The crossings of two bands from this four just produce the pair of the Weyl points. Due to this complexity of the charge-carriers spectrum, numerical calculations of the total electron-band structure of Weyl semimetals are commonly used to find the Fermi surface and its characteristics in these materials [8, 9, 10, 11, 12].
@@
However, Weng et al. [7] indicated the general form of the Hamiltonian that describes the four bands originating from the band-contact line lying in the mirror-reflection plane (the plane is also assumed to contain the two-fold or four-fold symmetry axis). Such planes just exist in semimetals of the TaAs family. With this general form, we formulate the four-bands model describing the spectrum of the charge carriers near the two close Weyl points in these materials (Sec. II). Within this model, we find conditions for the existence of the Weyl points and explicit analytical formulas for $\varepsilon^{W}$ and for the dispersion of the four bands along the line connecting the close Weyl points in the Brillouin zone. It also follows from the results of this section that the four bands and the cross-sectional areas of the Fermi-surface can be found by numerically solving a quartic equation, avoiding the complicated band-structure calculations. In other words, the model makes it possible to simply analyze various experimental data obtained with oscillation effects. In Sec. III, using NbP as an example, we demonstrate how the parameters of the model can be found for the so-called [1] W1 and W2 Weyl points, comparing the characteristic features of the model spectrum with the data of the band-structure calculations for this semimetal [8]. In Sec. IV, the obtained results are discussed, and conclusions are presented in Sec. V. Appendixes contain some mathematical details.
@@ II Charge-carriers spectra near Weyl points
Figure 1: The cross section (hexagon) of the first Brillouin zone of NbP by the mirror-reflection plane $p_{y}=0$, and the band-contact rings in this plane. The red rhombi depict the projections of the Weyl points onto this plane. The upper right ring (enclosed by the dashed square) is also shown in an enlarged scale. For this ring, the W1 and W2 Weyl points are marked, near which the energy bands were calculated in Ref. [8] (see Figs. 2, 3, 5-8, and Tables 1, 3). For these Weyl points, the angle $\theta$ between the $p_{z}$ axis and the $p_{\|}$ direction is indicated.
@@
Figure 2: Dispersion of the four bands along the $p_{y}$ direction ($k_{y}=p_{y}/\hbar$) at $p_{x}=p_{x}^{W}$, $p_{z}=p_{z}^{W}$, Eq. (21). The point ($p_{x}^{W}$$p_{z}^{W}$) in the mirror-reflection plane $p_{y}=0$ is the projection of the two Weyl points onto this plane. The bands are plotted for $m_{3}=m_{5}=0$ and the values of the other parameters presented in Table 2. (Interestingly, the same bands are obtained if the values of $m_{2}$ and $m_{4}/m_{6}$ are replaced by $m_{2}=10$ meV and $m_{4}/m_{6}=0.578$). The red cycles are the data of Fig. 8a in Ref. [8] for the W1 point in NbP (see Fig. 1) ; $\varepsilon_{i}$ mark the energy bands. All the energies are measured from the Fermi level (red dashed line).
@@ III Charge-carrier spectrum for $\rm NbP$}
III.1 The W1 points
Table 1: Quantities characterizing the W1 points in NbP. The values of these quantities were obtained from numerical calculations [8] of the electron-band structure of NbP near the W1 point marked in Fig. 1. All the energies are reckoned from the Fermi level, $\varepsilon_{i}(0)$ are the energies of the bands at the point $(p_{x}^{W1},0,2\pi\hbar/c)$ of the mirror-reflection plane $p_{y}=0$ (only $\varepsilon_{i}(0)$ for $i=2,3,4$ can be extracted from the data of Ref. [8]). Here $a=3.334$ Å.
@@ III.1 The W1 points
Figure 5: Dispersion of the two crossing bands along the $p_{z}$ direction ($k_{z}=p_{z}/\hbar$) in NbP at $p_{x}=p_{x}^{W1}$, $p_{y}=p_{y}^{W1}$, Eq. (29). The black solid lines show the bands for the parameters presented in Table 2 and for $R=0.2(2\pi\hbar/a)$, $d^{2}\varepsilon_{0}/dp_{\|}^{2}\approx 0.37/m$, and $v_{\|,\|}=0.79\times 10^{4}$ m/s (which corresponds to $v_{\|}=0.95\times 10^{4}$ m/s) where $a=3.334$ Å and $m$ is the free-electron mass. Note that the black solid lines remain practically unchanged if $v_{\|,\|}<10^{4}$ m/s. The red cycles are the data of Fig. 8c in Ref. [8]. These data were obtained for the W1 point marked in Fig. 1. All the energies are measured from the Fermi level (red dashed line).
@@ III.2 The W2 points
Consider a pair of the close W2 points in NbP (this pair is marked in Fig. 1). Condition (14) means that $m_{4}=-(m_{4}/m_{6})m_{5}$ near these points. Like for the W1 points, the dispersion of the bands along the $p_{y}$ axis, Eq. (21), is determined by the five parameters: $t_{2}$, $\bar{\varepsilon}^{W2}$, $m_{2}$, $m_{4}/m_{6}$, and $\tilde{m}\equiv\sqrt{m_{1}^{2}+m_{5}^{2}+m_{6}^{2}}$. Using the values of the quantities presented in Table 3, these parameters are found (Table 4), see also Fig. 6. As in the case of the W1 points, we find two sets of the parameters for the data of Table 3. The first set is obtained under condition (20), whereas the second set is derived at the opposite inequality. Both the sets describe the dispersion of the bands along the $p_{y}$ direction equally well. As in the case of the W1 points, the second set of the parameters leads to the appearance of additional Weyl points in the mirror-reflection plane. Such points have never been detected in the band-structure calculations and experiments, and so we exclude the second sets of the parameters from our consideration.
@@ IV Discussion
The described origin of the Weyl points in the topological materials of the TaAs family points out to a possible direction for searching for other noncentrosymmetric Weyl semimetals. It is well known [22] that in neglect of the spin-orbit interaction, an accidental contact of two bands of the opposite parities can occur along lines in the mirror-reflection planes of noncentrosymmetric crystals. For example, the inversion of these bands at some point of the plane is sufficient for the appearance of such a line. Then, the Hamiltonian of Sec. II can be applied to describe the energy spectrum near this nodal line. If the plane contains a two, four, or six-fold rotation axis, which may be denote as $z$ axis, and the nodal line in this plane crosses the plane, $p_{z}=0$, or the Brillouin-zone faces perpendicular to $p_{z}$, the conditions like Eqs. (II) can lead to existence of the W1 points in such a material.
@@
where $d_{z}=a^{\prime}p_{\perp}$ is described by formula (8), and $\tilde{m}\equiv\sqrt{m_{1}^{2}+m_{5}^{2}+m_{6}^{2}}$. Note that at $m_{1}\neq 0$, Eq. (32) gives the band-contact line, $d_{z}=a^{\prime}p_{\perp}=(m_{4}m_{6}-m_{3}m_{5})/m_{1}$, in the plane $p_{y}=0$. Along this line, one has $\epsilon_{s,1}=\epsilon_{s,-1}$ for $s=\pm 1$. The appearance of this line is due to the fail of condition (20).
@@ V Conclusions
Appendix A Energy spectrum in the immediate vicinity of a Weyl point.
@@
Note that for the direction along $p_{\|}$, formulas (35), (38), (42) contain not only the linear terms in $p_{\|}$ but also the quadratic terms in this component of the quasimomentum. These quadratic terms can be significant because the velocities $v_{1}$, $_{2}, v~\tilde{v}$, $_{\perp}, vv_{\|}$ associated with the spin-orbit interaction are relatively small in the TaAs family of the semimetals [due to formula (18), the term $(p_{y}^{W})^{2}t_{2}t_{2}^{\prime}$ has the same order of magnitude as the other terms in $c_{0}$ defined by the first expression in (A)]. However, if a direction, along which the bands are considered, noticeably deviates from the $p_{\|}$ axis, the quadratic terms in $\Delta d_{z}$ and $\Delta\varepsilon$ can be omitted. In this situation, it is convenient to rewrite formulas (38), (42) in terms of the variables $\Delta p_{x}$, $\Delta p_{y}$ $\Delta p_{z}$. This representation is useful when the bands are analyzed along the $p_{x}$, $p_{y}$ axes for the W1 points and along the $p_{x}$, $p_{y}$, $p_{z}$ directions in the case of the W2 points.
@@
The use of Eqs. (II)–(27) with conditions (28) and formula (46) lead to the minimal model of the four-bands spectrum (Sec. II) that reproduces the main feature of the Weyl points, the linear splitting of the bands in all the directions of the quasimomentum $\bf p$} at ${\bf p}\to{\bf p}^{W}$.
@@
$(a^{\prime})^{2}$ $=$ $\frac{q_{1}^{2}}{q_{2}}\frac{Q_{xx}-Q_{zz}\tan^{2}\theta}{1-\tan^{2}\theta},$ (47)
@@
where $Q_{xx}=(v_{x+}^{W2}-v_{x-}^{W2})^{2}/4$, $Q_{xx}=(v_{z+}^{W2}-v_{z-}^{W2})^{2}/4$, and $q_{1}$, $q_{2}$ are defined by Eqs. (39). The formula for $v_{\|,\|}^{2}$ shows that in the minimal model, there is a critical value $\theta=\theta_{0}$,
@@
at which the velocity $v_{\|,\|}$ becomes zero. On the other hand, at $|\theta|<|\theta_{0}|$, $v_{\|,\|}$ sharply increases, $v_{\|,\|}\propto\sqrt{|\theta_{0}|-|\theta|}$. For example, for $v_{x\pm}$, $v_{z\pm}$ from Table 3, we find $|\theta_{0}|\approx 128^{\circ}$, whereas at $\theta=125^{\circ}$, $v_{\|,\|}$ already reaches $1.08\times 10^{5}$ m/s. Thus, a closeness of $\theta$ to $\theta_{0}$ is the necessary condition for the applicability of the minimal model to the description of the four-bands spectrum.
@@ Appendix C Cross-sectional areas of Fermi surface near Weyl points
Figure 9: The cross sections of the Fermi surface for the third (outer curve) and forth (inner curve) bands of NbP by the plane perpendicular to the $p_{z}$ axis and passing through the W1 points (red rhombi). Here $k_{x}=p_{x}/\hbar$, $k_{y}=p_{y}/\hbar$, which are measured from the middle of the line segment connecting the Weyl points. The Fermi energy is shifted by $24$ meV downwards compared to its value in Table 2 and Fig. 2 in order to explain the notations $k_{y}^{(1a)}(k_{x}^{\prime})$, $k_{y}^{(1b)}(k_{x}^{\prime})$, $k_{y}^{(2)}(k_{x}^{\prime})$ introduced in the text $k_{x}^{\prime}$ coincides with $k_{x}$ for the plane perpendicular to $k_{z}$]. The upper segment D-E of the closed curve shows nonzero values of $k_{y}^{(1b)}(k_{x})$, whereas the segments B-A and B-C depict the functions $k_{y}^{(1a)}(k_{x})$ and $k_{y}^{(2)}(k_{x})$, respectively, in the regions where their imaginary parts vanish.

### v2

# Model for charge carrier spectra in topological semimetals of the TaAs family

@@ Abstract
We propose a four-band model describing the electron energy spectra near the Weyl points in the topological semimetals of the TaAs family (TaAs, TaP, NbAs, NbP). This model takes into account the fact that these Weyl points result from the band-contact lines which would exist in the mirror-reflection planes of these materials if the spin-orbit interaction were absent in them. Within this model, we obtain conditions for the existence of the Weyl points, determine their positions in the Brillouin zone, and derive the explicit formula for dispersion of the bands along the straight line connecting the two close Weyl points with opposite topological charges. Using NbP as an example, the values of the parameters defining the model spectrum are found. The obtained results show that for the semimetals of the TaAs family, the charge-carriers spectrum in the vicinity of the two close Weyl points can be analyzed without complex band-structure calculations.
@@ I Introduction
where $p_{i}$ ($i=x,y,z$) are the components of the quasimomentum, the velocities $v_{i}$ specify the splitting of the bands at the Weyl point ${\bf p}=0$ that has the energy $\varepsilon^{W}$, and the constant vector ${\bf a}=(a_{x},a_{y},a_{z})$ with the dimension of the velocity determines the so-called tilt of the spectrum. Moreover, the use of the additional simplification ${\bf a}=0$ is not rare in publications. However, only pairs of Weyl points with the opposite topological charges (the Chern numbers) can arise in the Brillouin zone of crystals, and the distance between these points of each pair is relatively small. This means that the region of the Brillouin zone where the linear dispersion of the bands really occurs is small, and a deviation from the linear dispersion can manifest itself even for a small difference between the Fermi energy $\varepsilon_{F}$ and the energy $\varepsilon^{W}$ of the Weyl points in the pair. Therefore, in order to correctly find the Fermi surface and its extremal cross sections, calculate the Landau levels and various physical quantities for the Weyl semimetals in a magnetic field, it may be necessary to consider the energy spectrum of the pair of the Weyl points [3] (see also recent papers [4, 5, 6, 7]). Moreover, if $|\varepsilon_{F}-\varepsilon^{W}|$ is larger than the energy barrier separating the points, for the correct description of the charge-carriers spectrum, it is necessary to take into account four close energy bands that exist near the Weyl points. These four bands result from the nodal line (lying in the mirror-reflection plane) which would exist in these semimetals if the spin-orbit were absent in them [8]. The crossings of two bands from this four just produce the pair of the Weyl points. Due to this complexity of the charge-carriers spectrum, numerical calculations of the total electron-band structure of Weyl semimetals are commonly used to find the Fermi surface and its characteristics in these materials [9, 10, 11, 12, 13].
@@
However, Weng et al. [8] indicated the general form of the Hamiltonian that describes the four bands originating from the band-contact line lying in the mirror-reflection plane (the plane is also assumed to contain the twofold or fourfold symmetry axis). Such planes just exist in semimetals of the TaAs family. With this general form, we formulate the four-band model describing the spectrum of the charge carriers near the two close Weyl points in these materials (Sec. II). Within this model, we find conditions for the existence of the Weyl points and explicit analytical formulas for $\varepsilon^{W}$ and for the dispersion of the four bands along the line connecting the close Weyl points in the Brillouin zone. It also follows from the results of this section that the four bands and the cross-sectional areas of the Fermi surface can be found by numerically solving a quartic equation, avoiding the complicated band-structure calculations. In other words, the model makes it possible to simply analyze various experimental data obtained with oscillation effects. In Sec. III, using NbP as an example, we demonstrate how the parameters of the model can be found for the so-called [1] W1 and W2 Weyl points, comparing the characteristic features of the model spectrum with the data of the band-structure calculations for this semimetal [9]. In Sec. IV, the obtained results are discussed, and conclusions are presented in Sec. V. Appendixes contain some mathematical details.
@@ II Charge-carriers spectra near Weyl points
Figure 1: Cross section (hexagon) of the first Brillouin zone of NbP by the mirror-reflection plane $p_{y}=0$, and the band-contact rings in this plane. The red rhombi depict the projections of the Weyl points onto this plane. The upper right ring (enclosed by the dashed square) is also shown in an enlarged scale. For this ring, the W1 and W2 Weyl points are marked, near which the energy bands were calculated in Ref. [9] (see subsequent Figs. 2, 3, 5-8, and Tables 1, 3). For these Weyl points, the angle $\theta$ between the $p_{z}$ axis and the $p_{\|}$ direction is indicated.
@@
Figure 2: Dispersion of the four bands along the $p_{y}$ direction ($k_{y}=p_{y}/\hbar$) at $p_{x}=p_{x}^{W}$, $p_{z}=p_{z}^{W}$, Eq. (21). The point ($p_{x}^{W}$$p_{z}^{W}$) in the mirror-reflection plane $p_{y}=0$ is the projection of the two Weyl points onto this plane. The bands are plotted for $m_{3}=m_{5}=0$ and the values of the other parameters presented in Table 2. (Interestingly, the same bands are obtained if the values of $m_{2}$ and $m_{4}/m_{6}$ are replaced by $m_{2}=10$ meV and $m_{4}/m_{6}=0.578$.) The red cycles are the data of Fig. 8a in Ref. [9] for the W1 point in NbP (see Fig. 1) ; $\varepsilon_{i}$ mark the energy bands. All the energies are measured from the Fermi level (red dashed line).
@@ III Charge-carrier spectrum for $\rm NbP$}
III.1 W1 points
Table 1: Quantities characterizing the W1 points in NbP. The values of these quantities were obtained from numerical calculations [9] of the electron-band structure of NbP near the W1 point marked in Fig. 1. All the energies are reckoned from the Fermi level, $\varepsilon_{i}(0)$ are the energies of the bands at the point $(p_{x}^{W1},0,2\pi\hbar/c)$ of the mirror-reflection plane $p_{y}=0$ (only $\varepsilon_{i}(0)$ for $i=2,3,4$ can be extracted from Fig. 8a of Ref. [9]). Here $a=3.334$ Å [9].
@@ III.1 W1 points
Figure 5: Dispersion of the two crossing bands along the $p_{z}$ direction ($k_{z}=p_{z}/\hbar$) in NbP at $p_{x}=p_{x}^{W1}$, $p_{y}=p_{y}^{W1}$, Eq. (29). The black solid lines show the bands for the parameters presented in Table 2 and for $R=0.2(2\pi\hbar/a)$, $d^{2}\varepsilon_{0}/dp_{\|}^{2}\approx 0.37/m$, and $v_{\|,\|}=0.79\times 10^{4}$ m/s (which corresponds to $v_{\|}=0.95\times 10^{4}$ m/s) where $a=3.334$ Å, and $m$ is the free-electron mass. Note that the black solid lines remain practically unchanged if $v_{\|,\|}<10^{4}$ m/s. The red cycles are the data of Fig. 8c in Ref. [9]. These data were obtained for the W1 point marked in Fig. 1. All the energies are measured from the Fermi level (red dashed line).
@@ III.2 W2 points
Consider a pair of the close W2 points in NbP (this pair is marked in Fig. 1). Condition (14) means that $m_{3}=-(m_{4}/m_{6})m_{5}$ near these points. Like for the W1 points, the dispersion of the bands along the $p_{y}$ axis, Eq. (21), is determined by the five parameters: $t_{2}$, $\bar{\varepsilon}^{W2}$, $m_{2}$, $m_{4}/m_{6}$, and $\tilde{m}\equiv\sqrt{m_{1}^{2}+m_{5}^{2}+m_{6}^{2}}$. Using the values of the quantities presented in Table 3, these parameters are found (Table 4), see also Fig. 6. As in the case of the W1 points, we find two sets of the parameters for the data of Table 3. The first set is obtained under condition (20), whereas the second set is derived at the opposite inequality. Both the sets describe the dispersion of the bands along the $p_{y}$ direction equally well. As in the case of the W1 points, the second set of the parameters leads to the appearance of additional Weyl points in the mirror-reflection plane. Such points have never been detected in the band-structure calculations and experiments, and so we exclude the second sets of the parameters from our consideration.
@@ IV Discussion
The described origin of the Weyl points in the topological materials of the TaAs family points out to a possible direction for searching for other noncentrosymmetric Weyl semimetals. It is well known [23] that in neglect of the spin-orbit interaction, an accidental contact of two bands of the opposite parities can occur along lines in the mirror-reflection planes of noncentrosymmetric crystals. For example, the inversion of these bands at some point of the plane is sufficient for the appearance of such a line. Then, the Hamiltonian of Sec. II can be applied to describe the energy spectrum near this nodal line. If the plane contains a two-, four-, or six-fold rotation axis, which may be denoted as the $z$ axis, and the nodal line in this plane crosses the plane, $p_{z}=0$, or the Brillouin-zone faces perpendicular to $p_{z}$, the conditions like Eqs. (II) can lead to existence of the W1 points in such a material.
@@
where $d_{z}=a^{\prime}p_{\perp}$ is described by formula (8), and $\tilde{m}\equiv\sqrt{m_{1}^{2}+m_{5}^{2}+m_{6}^{2}}$. Note that at $m_{1}\neq 0$, Eq. (32) gives the band-contact line, $d_{z}=a^{\prime}p_{\perp}=(m_{4}m_{6}-m_{3}m_{5})/m_{1}$, in the plane $p_{y}=0$. Along this line, one has $\epsilon_{s,1}=\epsilon_{s,-1}$ for $s=\pm 1$. The appearance of these lines is due to the fail of condition (20).
@@ V Conclusions
Appendix A Energy spectrum in the immediate vicinity of a Weyl point
@@
Note that for the direction along $p_{\|}$, formulas (35), (38), (42) contain not only the linear terms in $p_{\|}$ but also the quadratic terms in this component of the quasimomentum. These quadratic terms can be significant because the velocities $v_{1}$, $_{2}, v~\tilde{v}$, $_{\perp}, vv_{\|}$ associated with the spin-orbit interaction are relatively small in the TaAs family of the semimetals [due to formula (18), the term $(p_{y}^{W})^{2}t_{2}t_{2}^{\prime}$ has the same order of magnitude as the other terms in $c_{0}$ defined by the first expression in Eqs. (A)]. However, if a direction, along which the bands are considered, noticeably deviates from the $p_{\|}$ axis, the quadratic terms in $\Delta d_{z}$ and $\Delta\varepsilon$ can be omitted. In this situation, it is convenient to rewrite formulas (38), (42) in terms of the variables $\Delta p_{x}$, $\Delta p_{y}$ $\Delta p_{z}$. This representation is useful when the bands are analyzed along the $p_{x}$, $p_{y}$ axes for the W1 points and along the $p_{x}$, $p_{y}$, $p_{z}$ directions in the case of the W2 points.
@@
The use of Eqs. (II)–(27) with conditions (28) and formula (46) lead to the minimal model of the four-band spectrum (Sec. II) that reproduces the main feature of the Weyl points, the linear splitting of the bands in all the directions of the quasimomentum $\bf p$} at ${\bf p}\to{\bf p}^{W}$.
@@
$(a^{\prime})^{2}$ $=$ $\frac{q_{1}^{2}}{q_{2}}\frac{(Q_{xx}-Q_{zz}\tan^{2}\theta)}{(1-\tan^{2}\theta)},$ (47)
@@
where $Q_{xx}=(v_{x+}^{W2}-v_{x-}^{W2})^{2}/4$, $Q_{zz}=(v_{z+}^{W2}-v_{z-}^{W2})^{2}/4$, and $q_{1}$, $q_{2}$ are defined by Eqs. (39). The formula for $v_{\|,\|}^{2}$ shows that in the minimal model, there is a critical value $\theta=\theta_{0}$,
@@
at which the velocity $v_{\|,\|}$ becomes zero. However, at $|\theta|<|\theta_{0}|$, $v_{\|,\|}$ sharply increases, $v_{\|,\|}\propto\sqrt{|\theta_{0}|-|\theta|}$. For example, for $v_{x\pm}$, $v_{z\pm}$ from Table 3, we find $|\theta_{0}|\approx 128^{\circ}$, whereas at $\theta=125^{\circ}$, $v_{\|,\|}$ already reaches $1.08\times 10^{5}$ m/s. Thus, a closeness of $\theta$ to $\theta_{0}$ is the necessary condition for the applicability of the minimal model to the description of the four-bands spectrum.
@@ Appendix C Cross-sectional areas of Fermi surface near Weyl points
Figure 9: Cross sections of the Fermi surface for the third (outer curve) and forth (inner curve) bands of NbP by the plane perpendicular to the $p_{z}$ axis and passing through the W1 points (red rhombi). Here $k_{x}=p_{x}/\hbar$, $k_{y}=p_{y}/\hbar$, which are measured from the middle of the line segment connecting the Weyl points. The Fermi energy is shifted by $24$ meV downwards compared to its value in Table 2 and Fig. 2 in order to explain the notations $k_{y}^{(1a)}(k_{x}^{\prime})$, $k_{y}^{(1b)}(k_{x}^{\prime})$, $k_{y}^{(2)}(k_{x}^{\prime})$ introduced in the text $k_{x}^{\prime}$ coincides with $k_{x}$ for the plane perpendicular to $k_{z}$]. The upper segment D-E of the closed curve shows nonzero values of $k_{y}^{(1b)}(k_{x})$, whereas the segments B-A and B-C depict the functions $k_{y}^{(1a)}(k_{x})$ and $k_{y}^{(2)}(k_{x})$, respectively, in the regions where their imaginary parts vanish.